# Time-Correlated Video Bridge Matching

### v1

@@ Appendix B ADDITIONAL QUANTITATIVE RESULTS
Appendix C INITIALIZATION EXPERIMENTS
In this section, we explore options for initializing or representing input data for bridge-based methods used in our work, namely for Bridge Matching with Browninan Bridge (BM) and Time-Correlated Video Bridge Matching (TCVBM). The interest in initialization is primarily due to the fact that presumably these methods are better suited to data-to-data translation tasks. The interest in exploring the effect of input data initialization on the quality of model performance stems primarily from the assumption that bridge-based approaches are better suited for data-to-data translation tasks.
C.1 Frame Interpolation
@@
C.2 Image-to-Video Generation
@@
C.3 Video Super Resolution
@@
Appendix D HYPERPARAMETERS SEARCHING
Here we compare different values of hyperparameters, namely the noise scaling value $\epsilon$ and the coefficient $\alpha$, which determines the degree of impact of the matrix $\mathbf{A}$ as $\widetilde{\mathbf{A}}:=\alpha\mathbf{A}$. Table 11 shows that in the case of frame interpolation, it is impossible to identify a clear dependence of the generation quality on the hyperparameters used, however, a sufficient amount of noise and not large values for the $\alpha$ coefficient are optimal. The results for video super resolution in Table 11 demonstrate that small values of $\epsilon$ and $\alpha$ are optimal for this task, which does not contradict the results for frame interpolation.
@@ Appendix D HYPERPARAMETERS SEARCHING
Appendix E DYNAMICAL CORRELATION
E.1 Theory
@@
$\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})=\mathbb{E}[\mathbf{X}_{t}|\mathbf{X}_{0}]=e^{\mathbf{A}F(t)}\mathbf{X}_{0}+\int_{0}^{t}e^{\mathbf{A}(F(t)-F(s))}f(s)\,\mathbf{b}\,ds.$
@@
$\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})=e^{\mathbf{A}F(t)}\mathbf{X}_{0}+\big(e^{\mathbf{A}F(t)}-I\big)\mathbf{A}^{-1}\mathbf{b}.$
@@
$\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})=e^{\mathbf{A}F(t)}\mathbf{X}_{0}+\left(e^{\mathbf{A}F(t)}-I\right)\mathbf{A}^{-1}\mathbf{b}.$
Conditional variance. Let $\mathbf{Z}_{t}:=\mathbf{X}_{t}-\boldsymbol{\mu}_{t|0}$ be the centered process. Subtracting the mean SDE from the original SDE yields
@@
Define the covariance $\boldsymbol{\Sigma}_{t|0}:=\mathbb{E}[\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}]$. Using Itô’s rule for $\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}$,
@@
$\frac{d}{dt}\boldsymbol{\Sigma}_{t|0}=f(t)\,\mathbf{A}\,\boldsymbol{\Sigma}_{t|0}+f(t)\,\boldsymbol{\Sigma}_{t|0}\mathbf{A}^{\top}+\epsilon\,I,\quad\boldsymbol{\Sigma}_{t|0}=0.$
@@
$\boldsymbol{\Sigma}_{t,t^{\prime}|0}:=\mathrm{Cov}(\mathbf{Z}_{t},\mathbf{Z}_{t^{\prime}}^{\top})=\epsilon\int_{0}^{t}e^{\mathbf{A}(F(t)-F(s))}\,e^{\mathbf{A}(F(t^{\prime})-F(s))^{\top}}\,ds.$
@@
$\boldsymbol{\Sigma}_{t,t^{\prime}|0}=e^{\mathbf{A}(F(t^{\prime})-F(t))}\,\epsilon\int_{0}^{t}e^{2\mathbf{A}(F(t)-F(s))}\,ds=e^{\mathbf{A}\,[F(t^{\prime})-F(t)]}\,\boldsymbol{\Sigma}_{t|0}.$
Summary. Thus, all three components: mean $\boldsymbol{\mu}_{t|0}(\mathbf{X}_{0})$, variance $\boldsymbol{\Sigma}_{t|0}$, and cross-covariance $\boldsymbol{\Sigma}_{t,t^{\prime}|0}$ are derived and can be used further in the same way as in the original case of $f(t)=1$.
E.2 Experimental Results
The continuous and time-decreasing function $f(t)$ sets the increasing values of the matrix $\mathbf{A}$ in the inverse diffusion process when $t\rightarrow 0$. Thus, the correlation of frames with each other in the generated video increases in the last steps of the inference. We conduct a series of experiments to investigate the effect of dynamic correlation and the choice of the function $f(t)$. We use the same experimental setting as described in Section 5.1 of the main paper, with a number of optimizer steps set to $120,000$. The frame interpolation results are presented in Table 12. As can be seen, our experiments do not demonstrate the advantages of using dynamic correlation compared to using constant values of the matrix $\mathbf{A}$. However, we observe significant differences in the quality of the results depending on the function $f(t)$. This demonstrates a complex structure of the relationship between video frames in the diffusion process, which our framework provides in the constant linear approximation.
@@ E.2 Experimental Results
Appendix F ADDITIONAL VISUAL EXAMPLES
F.1 Frame Interpolation
@@
F.2 Image-to-Video Generation
@@
F.3 Video Super Resolution

### v2

@@ Appendix C Additional Quantitative Results
Appendix D Initialization Experiments
In this section, we explore options for initializing or representing input data for bridge-based methods used in our work, namely for Bridge Matching with Browninan Bridge (BM) and Time-Correlated Video Bridge Matching (TCVBM). The interest in exploring the effect of input data initialization on the quality of model performance stems primarily from the assumption that bridge-based approaches are better suited for data-to-data translation tasks.
D.1 Frame Interpolation
@@
D.2 Image-to-Video Generation
@@
D.3 Video Super Resolution
@@
Appendix E Hyperparameters Searching
Here we compare different values of hyperparameters, namely the noise scaling value $\epsilon$ and the coefficient $\alpha$, which determines the degree of impact of the matrix $\mathbf{A}$ as $\widetilde{\mathbf{A}}:=\alpha\mathbf{A}$. Tables 10 and  11 show that in the case of frame interpolation and image-to-video generation, it is impossible to identify a clear dependence of the generation quality on the hyperparameters used, however, a sufficient amount of noise and not large values for the $\alpha$ coefficient are optimal. The results for video super resolution in Table 12 demonstrate that small values of $\epsilon$ and $\alpha$ are optimal for this task, which does not contradict the results for frame interpolation.
@@ Appendix E Hyperparameters Searching
Table 11: The results of TCVBM training with various hyperparameters $\epsilon$ and $\alpha$ for image-to-video generation. The best values in column are bold, second best values are underlined.
0.1 0.1 55.46 0.2670 10.80 0.587
0.1 1 51.62 0.2578 10.74 0.583
0.1 10 59.57 0.2571 10.67 0.585
1 0.1 51.80 0.2615 10.70 0.590
1 1 44.96 0.2575 10.75 0.591
1 10 51.27 0.2613 10.60 0.583
10 0.1 44.90 0.2610 10.67 0.588
10 1 44.22 0.2584 10.70 0.591
10 10 44.58 0.2597 10.58 0.583
@@
Appendix F Computational Cost Analysis
As described in Algorithm 2, inference consists of two alternating steps: (i) computing $\widehat{\mathbf{X}}_{0}^{\phi}(\mathbf{X}_{t_{n}},t_{n})$ and (ii) sampling from the Gaussian distribution given in Eq. 9. Sampling can be done using reparameterization:
$\mathbf{X}_{t^{\prime}}=\mathbf{\mu}_{t|0,t^{\prime}}(\mathbf{X}_{0})+(\mathbf{\Sigma}_{t|0,t^{\prime}}^{1/2})^{\top}\mathbf{Z},\quad\mathbf{Z}\sim\mathcal{N}(0,I),$
which requires the evaluation of the mean and covariance:
$\mathbf{\mu}_{t|0,t^{\prime}}=\mathbf{\mu}_{t|0}(\mathbf{X}_{0})+\mathbf{\Sigma}_{t|0}\mathbf{\Sigma}_{t^{\prime}|0}^{-1}(\mathbf{X}_{t^{\prime}}-\mathbf{\mu}_{t^{\prime}|0}(\mathbf{X}_{0})),\quad$
$\mathbf{\Sigma}_{t|0,t^{\prime}}=\mathbf{\Sigma}_{t|0}-\mathbf{\Sigma}_{t|0}\mathbf{\Sigma}_{t^{\prime}|0}^{-1}\mathbf{\Sigma}_{t|0}.$
In turn,
$\mathbf{\mu}_{t|0}(\mathbf{X}_{0})=e^{\mathbf{A}t}\mathbf{X}_{0}+\left(e^{\mathbf{A}t}-I\right)\mathbf{A}^{-1}\mathbf{b},\quad\mathbf{\Sigma}_{t|0}=\epsilon\frac{e^{2\mathbf{A}t}-I}{2}\mathbf{A}^{-1}.$
In total, almost all these operations can be cached except for 5 matrix multiplications: $(\mathbf{\Sigma}_{t|0,t^{\prime}}^{1/2})^{\top}\mathbf{Z}$, $e^{\mathbf{A}t}\mathbf{X_{0}}$, $(\mathbf{\Sigma}_{t|0}\mathbf{\Sigma}_{t^{\prime}|0}^{-1})\mathbf{X}_{t^{\prime}}$, $e^{\mathbf{A}t^{\prime}}\mathbf{X_{0}}$, $(\mathbf{\Sigma}_{t|0}\mathbf{\Sigma}_{t^{\prime}|0}^{-1})(e^{\mathbf{A}t^{\prime}}\mathbf{X_{0}})$. All these multiplications are for tensors of size $F\times F$ ($F$ is the number of frames in the video) with a tensor of size $F\times C\times H\times W$, where $C$ is the number of channels, $H$ is the height, and $W$ is the width of the video. This requires $O(F^{2}CHW)$ operations. For comparison, a single convolutional layer with $C$ input channels, $C_{\text{out}}$ output channels, and kernel size $K\times K$ applied to all $F$ frames requires $O(FCHWK^{2}C_{\text{out}})$ operations. Since typically $F\sim 10{-}100$, $K^{2}\sim 10$, and $C_{\text{out}}\sim 10$, we have
$\frac{F^{2}CHW}{FCHWK^{2}C_{\text{out}}}=\frac{F}{K^{2}C_{\text{out}}}\sim 0.1-1.$
Thus, each of these 5 operations is comparable to requiring less computation than 1 convolution layer. Since a neural network requires many convolutional layers and additional nonlinear processing, the resulting overhead of our bridge update is comparable to only a few convolutional layers and therefore remains relatively small in practice.
Appendix G Dynamical Correlation
G.1 Theory
@@
$\bm{\mu}_{t|0}(\mathbf{X}_{0})=\mathbb{E}[\mathbf{X}_{t}|\mathbf{X}_{0}]=e^{\mathbf{A}F(t)}\mathbf{X}_{0}+\int_{0}^{t}e^{\mathbf{A}(F(t)-F(s))}f(s)\,\mathbf{b}\,ds.$
@@
$\bm{\mu}_{t|0}(\mathbf{X}_{0})=e^{\mathbf{A}F(t)}\mathbf{X}_{0}+\big(e^{\mathbf{A}F(t)}-I\big)\mathbf{A}^{-1}\mathbf{b}.$
@@
$\bm{\mu}_{t|0}(\mathbf{X}_{0})=e^{\mathbf{A}F(t)}\mathbf{X}_{0}+\left(e^{\mathbf{A}F(t)}-I\right)\mathbf{A}^{-1}\mathbf{b}.$
Conditional variance. Let $\mathbf{Z}_{t}:=\mathbf{X}_{t}-\bm{\mu}_{t|0}$ be the centered process. Subtracting the mean SDE from the original SDE yields
@@
Define the covariance $\bm{\Sigma}_{t|0}:=\mathbb{E}[\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}]$. Using Itô’s rule for $\mathbf{Z}_{t}\mathbf{Z}_{t}^{\top}$,
@@
$\frac{d}{dt}\bm{\Sigma}_{t|0}=f(t)\,\mathbf{A}\,\bm{\Sigma}_{t|0}+f(t)\,\bm{\Sigma}_{t|0}\mathbf{A}^{\top}+\epsilon\,I,\quad\bm{\Sigma}_{t|0}=0.$
@@
$\bm{\Sigma}_{t,t^{\prime}|0}:=\mathrm{Cov}(\mathbf{Z}_{t},\mathbf{Z}_{t^{\prime}}^{\top})=\epsilon\int_{0}^{t}e^{\mathbf{A}(F(t)-F(s))}\,e^{\mathbf{A}(F(t^{\prime})-F(s))^{\top}}\,ds.$
@@
$\bm{\Sigma}_{t,t^{\prime}|0}=e^{\mathbf{A}(F(t^{\prime})-F(t))}\,\epsilon\int_{0}^{t}e^{2\mathbf{A}(F(t)-F(s))}\,ds=e^{\mathbf{A}\,[F(t^{\prime})-F(t)]}\,\bm{\Sigma}_{t|0}.$
Summary. Thus, all three components: mean $\bm{\mu}_{t|0}(\mathbf{X}_{0})$, variance $\bm{\Sigma}_{t|0}$, and cross-covariance $\bm{\Sigma}_{t,t^{\prime}|0}$ are derived and can be used further in the same way as in the original case of $f(t)=1$.
G.2 Experimental Results
The continuous and time-decreasing function $f(t)$ sets the increasing values of the matrix $\mathbf{A}$ in the inverse diffusion process when $t\rightarrow 0$. Thus, the correlation of frames with each other in the generated video increases in the last steps of the inference. We conduct a series of experiments to investigate the effect of dynamic correlation and the choice of the function $f(t)$. We use the same experimental setting as described in Appendix B of the main paper, with a number of optimizer steps set to $120,000$. The frame interpolation results are presented in Table 13. As can be seen, our experiments do not demonstrate the advantages of using dynamic correlation compared to using constant values of the matrix $\mathbf{A}$. However, we observe significant differences in the quality of the results depending on the function $f(t)$. This demonstrates a complex structure of the relationship between video frames in the diffusion process, which our framework provides in the constant linear approximation.
@@ G.2 Experimental Results
Appendix H Additional Examples
H.1 Frame Interpolation
@@
H.2 Image-to-Video Generation
@@
H.3 Video Super Resolution